OpenStudy (anonymous):

find the 3rd order derivative of y=(3x+5)^3

OpenStudy (anonymous):

Taking the first order derivative => 3*(3x+5)^2 * 3 Taking the second order derivative => 3*2*(3x+5) * 3 * 3 = 54(3x+5) Taking the third order derivative => 54*3 = 162 Is that clear?

OpenStudy (anonymous):

Every time you take the derivative of y with respect to x we substitute (3x+5) for u and get u^2 but because u isn't a constant according to the chain rule we have to multiply with the derivative of u (=3) each time.

OpenStudy (anonymous):

* are all multiplications X right?

OpenStudy (anonymous):

Yes, a*b means a multiplied with b.